# Disentangling the Potential Impacts of Papers into Diffusion, Conformity, and Contribution Values

### v3

@@ 4.1. Datasets
To select samples for testing, we target papers published before the test observation point (see Figure 1 (b)). These papers should be written in English, have complete metadata, and include at least one reference, without imposing constraints on the number of citations received, to maintain the integrity of the data distribution. From the eligible pool, we randomly sampled 300,000 papers to compose the test set. Additionally, we gather all the associated metadata for these papers, encompassing titles, abstracts, authors, venues, and publication times. To imitate the practical scenario, the training observation point is set 5 years before the test observation point (equal to the observation time window $T$ and predicted interval $\Delta$) for non-overlapping. The validation observation point precedes the test observation point by 3 years. It is crucial to underscore that though predicted samples in the training set reappear in the test set, their input graphs and predicted values undergo significant variations due to dynamic contextual changes. This intrinsic challenge renders the task both complex and critical. Moreover, we meticulously categorize the papers in the results for thorough and detailed analysis.
4.2. Baseline Models
@@
    MUCas (Chen et al., 2022) first uses time-aware sampling to extract dynamic graph snapshots and generate node embeddings from the graph structure. Then, it utilizes MUG-caps, a hierarchical capsule network, to encode snapshots in different levels (order, node, and graph) and aggregate the final representations with the attention mechanism.
Dynamic GNN models: We employ recent dynamic GNN models to extract paper representations from the dynamic graphs collected the same as our model, taking the potential impact prediction as a graph regression task. Similar to HINTS, we adapt these models to concentrate on target-centric subgraphs rather than the entire graph.
@@
4.3. Implementation Details
All baselines and our proposed model are implemented with PyTorch and DGL/PyG. They are trained on an NVIDIA A800 80GB GPU and optimized with the Adam optimizer (Kingma and Ba, 2014). As for the baselines, we implement them directly using or referring to the official source codes. We develop the models with the training and validation sets and choose the final model with the best main metric on the validation set. Each dataset is divided into previous papers (included in the training set), fresh papers (new additions in the test set), and immediate papers (published at the test time point) for detailed analysis. We adhere to the recommended settings for all baselines. In our model across all datasets, the hop order $k$ is set to 2, corresponding top limits $K_{1}$ and $K_{2}$ are set to 100 and 20, the learning rate is set to 1e-4, the number of layers $L$ is set to 4, the co-cited/citing weight $\lambda$ is set to 0.5, and the categories of equal frequency bins $M$ are set to 5. Additionally, the batch size is set to 64 for CS and CHM, and 32 for PSY, the disentanglement weight $\beta$ is set to 0.5 for CS and PSY, and 1.0 for CHM, and the hidden dimension $d^{h}$ is set to 128 for CS and PSY, and 192 for CHM.
4.4. Evaluation Metric
@@
5. Results
5.1. Performance Comparison (RQ1)
To answer RQ1, we evaluate the performance of DPPDCC and other baselines. The comparison results are reported in Table 2 (Computer Science), 3 (Chemistry), and 4 (Psychology). We have the following observations:
Table 2. Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Computer Science. We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total. The best results are shown in bold, and the second results are underlined. Significant improvements over best baseline results are marked with .
CS model all prev fresh immed
[…]
C&C SciBERT 0.6847 0.3558 0.6077 0.4424 0.8054 0.1108 0.7860 0.0843
HINTS 0.6799 0.2773 0.6429 0.2028 0.7378 0.2268 0.7387 0.1718
MUCas 0.5079 0.6063 0.4331 0.6757 0.6250 0.4352 0.8261 -0.1079
Dynamic EGCN 0.8157 0.1174 0.8008 0.0531 0.8392 0.0210 0.8387 -0.1460
Dysat 0.6248 0.4355 0.5489 0.5029 0.7437 0.2321 0.8073 0.0089
ROLAND 0.5814 0.5245 0.5166 0.5634 0.6828 0.3764 0.7157 0.2511
Disen DisenGCN 0.8215 0.1428 0.7816 0.1022 0.8841 0.0206 0.8485 0.0236
DisenHAN 0.8138 0.0054 0.7411 0.1003 0.9276 -0.3220 1.0686 -0.8032
CAL 0.7323 0.2298 0.6859 0.2264 0.8050 0.0768 0.7940 0.0427
DisC 0.6687 0.3434 0.6144 0.3908 0.7536 0.1472 0.7900 -0.0735
DIDA 0.5941 0.4703 0.5018 0.5642 0.7387 0.2391 0.8181 -0.0207
ours DPPDCC 0.4473 0.6871 0.3783 0.7236 0.5553 0.5754 0.6035 0.4512
#improve (%) 11.93 13.32 12.65 7.09 11.15 32.22 15.68 79.67
Table 3. Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Chemistry. We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total. The best results are shown in bold, and the second results are underlined. Significant improvements over best baseline results are marked with .
CHM model all prev fresh immed
[…]
C&C SciBERT 0.5451 0.4648 0.4987 0.4709 0.6903 0.1477 0.6452 0.1653
HINTS 0.6024 0.3795 0.5848 0.3119 0.6574 0.2253 0.6449 0.1633
MUCas 0.4691 0.6168 0.4314 0.6306 0.5869 0.3624 0.8073 -0.3479
Dynamic EGCN 0.6424 0.2684 0.6177 0.1913 0.7196 0.0791 0.6976 0.0035
Dysat 0.5442 0.4573 0.5029 0.4411 0.6733 0.1998 0.6846 0.0582
ROLAND 0.5001 0.5410 0.4592 0.5374 0.6280 0.2942 0.6406 0.1800
Disen DisenGCN 0.6405 0.2735 0.6079 0.2151 0.7427 0.0332 0.6987 0.0387
DisenHAN 0.6132 0.3793 0.5817 0.3683 0.7118 0.0630 0.7313 -0.1466
CAL 0.5864 0.3928 0.5541 0.3707 0.6873 0.1158 0.6535 0.1135
DisC 0.5393 0.4770 0.5037 0.4578 0.6505 0.2387 0.6409 0.1719
DIDA 0.4978 0.5760 0.4642 0.5689 0.6028 0.3587 0.6426 0.1669
ours DPPDCC 0.4355 0.6503 0.4071 0.6329 0.5243 0.5042 0.5802 0.3247
#improve (%) 7.16 5.44 5.64 0.36 10.67 39.13 9.42 80.39
Table 4. Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Psychology. We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total. The best results are shown in bold, and the second results are underlined. Significant improvements over best baseline results are marked with .
PSY model all prev fresh immed
[…]
C&C SciBERT 0.6395 0.4308 0.6025 0.5041 0.7402 0.0068 0.7116 -0.1095
HINTS 0.7749 0.1988 0.7799 0.2007 0.7614 -0.0259 0.7064 -0.0804
MUCas 0.5207 0.6175 0.4804 0.6846 0.6306 0.2669 0.7708 -0.3521
Dynamic EGCN 0.7424 0.2305 0.7375 0.2469 0.7560 -0.0388 0.7102 -0.1386
Dysat 0.6861 0.2998 0.6607 0.3342 0.7555 -0.0170 0.7439 -0.2377
ROLAND 0.5643 0.5540 0.4996 0.6601 0.7406 0.0429 0.7926 -0.3765
Disen DisenGCN 0.6557 0.3945 0.6316 0.4364 0.7213 0.0758 0.7081 -0.0846
DisenHAN 0.7763 0.2455 0.8077 0.2187 0.6906 0.1387 0.6476 0.0522
CAL 0.5631 0.5646 0.5469 0.5975 0.6073 0.3256 0.6285 0.1166
DisC 0.5415 0.5915 0.5159 0.6343 0.6113 0.3233 0.6369 0.0963
DIDA 0.5878 0.5118 0.5604 0.5620 0.6624 0.1950 0.6681 -0.0092
ours DPPDCC 0.4463 0.7037 0.4138 0.7454 0.5349 0.4701 0.5633 0.2733
#improve (%) 14.29 13.96 13.86 8.87 11.92 44.38 10.38 134.40
(1) All models perform well on previous papers presented in the training set but experience degradation when dealing with fresh and immediate papers in the test set, particularly in terms of the LogR2. Many models fail to figure out the overall trend of immediate papers, which results in negative values in LogR2. Even the best baseline, MUCas, struggles with immediate papers due to limited graph structures, as it is a cascade prediction model designed for papers with sufficient citations. However, dynamic GNN models can effectively encode multi-hop information to complement the lack of cascade structures. Disentangled GNNs, in particular, effectively uncover crucial insights within both multi-hop graph structures and node representations, notably beneficial for datasets like Psychology, which are denser. Moreover, models such as SciBERT and HINTS leverage prior knowledge, like content information or citation distribution, to tackle these challenges. Comparatively, while SciBERT outperforms HINTS in fitting the increments of previous papers (SciBERT ¿ HINTS), it faces greater difficulty with fresh papers (SciBERT ¡ HINTS). This implies that graph structures exhibit better generality and transferability to unseen samples, while shared content information remains robust for immediate papers with limited graph structures, both alleviating issues related to data distribution shifts.
(2) Both dynamic and heterogeneous graphs are important for our task. On one side, since temporal information within dynamic graphs exposes the evolved trend, it plays a crucial role. Dynamic GNNs possess comparable performance across all datasets for almost all categories of papers. In particular, DIDA, which considers dynamic graphs, stands out among the disentangled GNNs, closely matching MUCas’ performance across all datasets. On the other side, leveraging heterogeneous information can further enhance the performance of dynamic GNNs, as demonstrated by ROLAND. This general dynamic GNN framework achieves competitive performance akin to MUCas by integrating heterogeneous information. However, proper modeling of heterogeneous information is crucial; otherwise, it might lead to performance degradation. For instance, DisenHAN, compared to its backbone DisenGCN, struggles to encode valuable information within heterogeneous graphs, notably within the complex dataset Psychology.
(3) While Disentanglement methods offer performance improvements, their adaptation to dynamic heterogeneous graphs is essential to achieve optimal flexibility. Notably, DIDA exhibits superior outcomes than ROLAND in LogR2 for Chemistry, even when relying solely on homogeneous graphs. Moreover, within the densest dataset, Psychology, Disentangled GNN models showcase impressive performance. Among them, CAL outperforms both MUCas and ROLAND in predicting fresh and immediate papers. However, in other scenarios, apart from DIDA, these methods face challenges in competing with task-specific models and dynamic GNNs.
(4) Our proposed model surpasses all baselines, by effectively modeling the dynamic heterogeneous graph and incorporating task-adjusted disentanglement. This leads to an average improvement of about 10% in MALE across all categories, along with a substantial advancement of over 30% in LogR2 for fresh and immediate papers. First, compared with traditional citation/cascade prediction models, it exhibits superior generalization and practicability through the fusion of content and context information within citation networks, as well as the disentanglement power. Content information aids in discerning the impact of previous papers and fortifies robustness towards immediate papers. Moreover, its adeptness in handling fresh and immediate papers stems from the rich information embedded within graph structures. Furthermore, we consider the unique properties of citation networks to encode complex high-order structural and temporal information, thereby outperforming common dynamic models. The proposed CGE facilitates the identification of influential papers related to the target paper based on comparative and multi-hop information. More importantly, our model introduces interpretability from disentanglement models without compromising performance and further adapts to the task for enhanced improvement. It excels in predicting precise citation increments and approaches the genuine contribution of papers by disentangling values derived from information diffusion and collective conformity, thus providing insights into the papers’ authentic impact.
5.2. Ablation Test (RQ2)
To answer RQ2, ablation tests are performed on both CGE and PDM, highlighting the significance of various modules. As demonstrated in Table 5, for CGE, (-co-cited/citing) removes the proposed co-cited/citing strengths and only uses the original attention scores, (-CompGAT) replaces the CompGAT with original GATv2 (Brody et al., 2022), and (-readout) simply obtains the snapshot representation through avg-pooling over all paper nodes without filtering. For PDM, (-diffusion) removes the diffusion disentanglement, (-conformity) omits the conformity disentanglement, (-orthog) disregards the orthogonal constraints, and (-disen) eliminates all perspectives and replaces with a single MLP for prediction.
Table 5. Results of ablation studies according to MALE.
@@ 5.2. Ablation Test (RQ2)
ours 0.4473 0.4355 0.4463
CGE -co-cited/citing 0.4650 0.5014 0.4528
-CompGAT 0.4722 0.4990 0.4582
-readout 0.5116 0.4404 0.4932
PDM -diffusion 0.4593 0.4530 0.4573
-conformity 0.5177 0.4747 0.4580
-orthog 0.4930 0.4388 0.4502
-disen 0.4606 0.4447 0.4670
We observe that: (1) All modules within DPPDCC play pivotal roles, as evidenced by the decrease in performance when any module is removed. Their significance varies across datasets due to diverse graph characteristics. (2) For CGE, the removal of type-specific attention readout may cause a notable performance drop. That could be attributed to the presence of excessive noisy information within multi-hop snapshots, which necessitates the filtering of aggregated information for better concentration. Specially, in chemistry, where subgraphs are sparser, simple avg-pooling might suffice to extract meaningful information. Furthermore, our proposed CompGAT effectively captures distinct comparative information, surpassing the original GATv2 in computer science and psychology. The incorporation of co-cited/citing strengths further enhances multi-hop information beyond the local neighborhood, resulting in improved performance. (3) Eliminating any disentanglement task adversely affects performance. Moreover, an evident decline is observed after removing the whole disentanglement module, especially in Psychology. Conformity rises as the paramount task among all disentanglement tasks. It focuses on the Matthew effect, pervasive and discernible in reality, thus significantly contributing to model performance.
5.3. Hyper-parameters Test (RQ3)
(a) $L$
(b) $\lambda$
(c) $\beta$
Figure 3. Results of hyper-parameters test in terms of MALE.
(a) $L$
(b) $\lambda$
(c) $\beta$
Figure 4. Results of hyper-parameters test in terms of LogR2.
To address RQ3, we conduct experiments to assess the influence of three critical hyper-parameters in our model design: the stacked layer number $L$, the importance of co-cited/citing strengths $\lambda$, and the weight of the disentanglement tasks $\beta$. To ensure robust experimental outcomes, we focus on tuning these specific hyper-parameters while keeping other hyper-parameters fixed at their optimal values. In the following section, we report the corresponding results when tuning other important hyper-parameters. The experiment results are described in Figure 3 and 4.
5.3.1. The Impact of Model Depth
The number of layers $L$ serves as a pivotal hyperparameter for GNN-based models. Within this study, we explore values of $L$ in the range of 2, 3, 4, 5. For GNN-based models, increasing the number of layers facilitates the capture of higher-order connectivity within the graph during information propagation and aggregation. However, it also introduces the challenge of over-smoothing, a trend we’ve observed. As the layer count $L$ increases, model performance typically exhibits an initial rise followed by a decline—a phenomenon attributed to the interplay between increased connectivity capturing and eventual over-smoothing effects. In DPPDCC, besides GNN’s inherent challenges, the incorporated multi-hop co-cited/citing information and the intricate interaction between time and space dimensions could contribute to this phenomenon. Achieving the optimal hyperparameter $L$ requires finding a delicate balance between the GNN’s depth and its multi-hop/multi-dimensional information, ensuring an equilibrium that maximizes performance.
5.3.2. The Impact of Co-cited/citing Strengths
The weight parameter of co-cited/citing strengths $\lambda$ governs the proportion of co-cited/citing strengths in the sum distributions relative to the original attention scores. Our study investigates optimal values of $\lambda$ within the range 0.25, 0.5, 0.75, 1.0. Setting $\lambda$ to 1.0 essentially disregards attention mechanisms, replacing them with fixed weights and reducing the model to a weighted GCN. Higher weights can potentially dominate significant signals learned by the model, while lower weights may inadequately incorporate additional information. Hence, a moderate value such as 0.5 might represent a suitable setting adaptable across various datasets, balancing the incorporation of additional information without overshadowing the model’s learned signals.
5.3.3. The Impact of Disentanglement Loss
The weight of disentanglement loss $\beta$ controls the importance of the disentanglement tasks relative to the main task. Our study explores optimal values of $\beta$ within the range 0.25, 0.5, 0.75, 1.0. Notably, the disentanglement weight $\beta$ exhibits greater stability in comparison to $L$ and $\lambda$, indicating consistent influence and effectiveness of the disentanglement task as an auxiliary task. Selecting an appropriate $\beta$ value can further enhance overall model performance, underscoring the significance of this parameter in refining the model’s effectiveness.
5.4. Visualization (RQ4)
To answer RQ4, we visualize both CGE and PDM in the test set of Chemistry to delve deeper into our model’s mechanisms. For CGE, we employ linear regressions to uncover relationships among the sum distribution $\alpha$, original attention distribution $e$, and co-cited/citing distribution $c$. This analysis aims to elucidate the interplay and dependencies among these distributions. For PDM, we first calculate correlations between predicted/ground-truth values and the perspective values/proportions. Next, we visualize the comprehensive compositions of disentangled values concerning time, category, and value. Subsequently, we delve into the representation distribution of conformity perspective to validate its rationality. This process offers insights into how different perspectives contribute to predictions.
5.4.1. Weighting Score Visualization
(6.48/16.74)

### v4

@@ D.1. Datasets
To select samples for testing, we target papers published before the test observation point. These papers should be written in English, have complete metadata, and include at least one reference without imposing constraints on the number of citations received to maintain the integrity of the data distribution. From the eligible pool, we randomly sampled 300,000 papers to compose the test set. Additionally, we gather all the associated metadata for these papers, encompassing titles, abstracts, authors, venues, and publication times. To imitate the practical scenario, the training observation point is set 5 years before the test observation point (equal to the observation time window $T$ and predicted interval $\Delta$) for non-overlapping. The validation observation point precedes the test observation point by 3 years. It is crucial to underscore that though predicted samples in the training set reappear in the test set, their input graphs and predicted values undergo significant variations due to dynamic contextual changes. This intrinsic challenge renders the task both complex and critical. Moreover, we meticulously categorize the papers in the results for thorough and detailed analysis.
Table 6. Detailed statistics of single input heterogeneous subgraph for DPPDCC.
dataset CS CHM PSY
phase train val test train val test train val test
nodes paper 204 222 257 149 187 254 361 401 460
author 446 501 613 505 646 907 1015 1173 1431
venue 72 80 92 35 41 50 92 104 123
time 23 24 26 22 24 27 32 34 37
edges cites 740 827 1016 480 662 1036 1607 1808 2085
writes 599 674 827 673 864 1218 1388 1598 1933
publishes 165 183 216 135 168 228 311 350 408
In Table 6, we present detailed statistics of the input data for DPPDCC to demonstrate the sufficiency of information within multi-hop heterogeneous subgraphs. For simplicity, we average the node and edge counts by time and sample. It is evident that the Psychology dataset is the densest, while the Computer Science and Chemistry datasets are comparable.
D.2. Descriptive Statistics
To validate the rationale behind the disentangled factors used in our work, we visualize the distribution of citation increments between papers potentially influenced by corresponding popularity factors and other normal papers. For the information diffusion effect, it is posited that documents cited by highly-cited papers are more likely to attract attention from other papers during dissemination, thereby accruing more citations. Conversely, for the collective conformity effect, it is argued that the cumulative citations received by a paper can reflect its partial reputation and consequently increase its likelihood of being cited. Highly-cited papers mentioned in both contexts are classified using a Pareto distribution method: papers providing over 80% of the total cumulative citations are deemed highly-cited. The papers cited by these highly-cited papers represent focused samples of the information diffusion effect, while the highly-cited papers themselves serve as focused samples of the collective conformity effect.
(1) In Figure 6, the significant deviation among different categories is evident, with papers cited by highly-cited ones generally receiving more citations. Notably, the CS dataset exhibits the most distinct pattern, with the left peak nearly disappearing, resulting in an overall distribution closer to a normal distribution. The model aims to extract these diffusion-related factors to model the information diffusion effect independently.
(2) Figure 7 illustrates the difference in attracting new citations between papers with accumulated citations and those with fewer citations, indicating that the collective conformity effect is more pronounced than the information diffusion effect. Notably, the distribution of focused samples in the CS dataset has essentially normalized, while other datasets also display significant peak differences. The model is expected to separately model the conformity effect, enabling it to focus more on aspects unrelated to existing cumulative citations and align more closely with contribution-related information.
(a) CS
(b) CHM
(c) PSY
Figure 6. Citation increment distribution of normal samples and focused samples cited by highly-cited papers for diffusion perspective.
(a) CS
(b) CHM
(c) PSY
Figure 7. Citation increment distribution of normal samples and focused highly-cited samples for conformity perspective.
D.3. Baseline Models
@@
    SPECTER2 (Cohan et al., 2020) generates document-level embeddings of scientific documents by pretraining a BERT-based model on relatedness signals extracted from the citation graph. Instead of fine-tuning the model’s adapters, we fully fine-tune the basic SPECTER2 for better performance.
@@
    CasSampling (Cheng et al., 2023) employs an importance-aware sampling method to extract critical components of cascade graphs based on outdegree, retaining the time series of the global graphs as node features. Subsequently, an attention aggregator is used to encode the global-level time flow.
    H2CGL (Cheng et al., 2023), the state-of-the-art citation prediction model, shares a similar dynamic setting and heterogeneous GNN framework with our approach. It encodes the dynamics of the citation network hierarchically and designs message-passing modules that consider the unique characteristics of citation networks. Additionally, it incorporates graph contrastive learning to refine representations, encouraging the model to encode potential citation information unrelated to specific topics.
Dynamic GNN models: We employ recent dynamic GNN models to extract paper representations from the dynamic graphs collected in the same way as our model, taking the potential impact prediction as a graph regression task. Similar to HINTS, we adapt these models to concentrate on target-centric subgraphs rather than the entire graph.
@@
    DyTed (Zhang et al., 2023a) designs a temporal-clips contrastive learning task along with a structural contrastive learning task to effectively disentangle time-invariant and time-varying representations. It further employs an adversarial learning framework with a disentanglement-aware discriminator for enhancement. EGCN and Dysat are utilized as the backbone dynamic GNNs of DyTed, with the best performance reported using Dysat.
    SILD (Zhang et al., 2023b) addresses distribution shifts in both dynamic and spectral domains. It builds upon DIDA with spectral transformation and proposes a disentangled spectrum mask to capture and utilize invariant and variant spectral patterns.
D.4. Implementation Details
All baselines and our proposed model are implemented using PyTorch and DGL (Wang et al., 2019)/PyG (Fey and Lenssen, 2019). They are trained on an NVIDIA A800 80GB GPU and optimized with the Adam optimizer (Kingma and Ba, 2014). For the baselines, we directly use or refer to the official source codes. Experiments are conducted with three random seeds, and the averaged results are reported. Models are developed using the training and validation sets, with the final model selected based on the best main metric on the validation set. Each dataset is divided into previous papers (included in the training set), fresh papers (new additions in the test set), and immediate papers (published at the test time point) for detailed analysis. We follow the recommended settings from the original papers or open-source codes for all baselines. In our model across all datasets, the hop order $k$ is set to 2, corresponding top limits $K_{1}$ and $K_{2}$ are set to 100 and 20, the learning rate is set to 1e-4, the number of layers $L$ is set to 4, the co-cited/citing weight $\lambda$ and the disentanglement weight $\beta$ are set to 0.5, and the categories of equal frequency bins $M$ are set to 5. Additionally, the batch size is set to 64 for CS and CHM, and 32 for PSY, and the hidden dimension $d$ is set to 128 for CS and PSY, and 192 for CHM.
D.5. Evaluation Metric
@@
Appendix E Complete Experimental Results
Table 7. The complete experimental results of total samples with standard deviations.
@@ Appendix E Complete Experimental Results
MALE↓ LogR2↑ MALE↓ LogR2↑ MALE↓ LogR2↑
SciBERT 0.6904±.007 0.3370±.016 0.5868±.030 0.3887±.054 0.6131±.019 0.4761±.034
SPECTER2 0.7482±.027 0.2337±.058 0.5832±.013 0.3906±.032 0.5935±.020 0.5082±.031
HINTS 0.9022±.108 -0.1227±.231 0.8265±.074 -0.1283±.160 0.8529±.081 -0.0198±.204
CasSampling 0.5061±.003 0.6112±.007 0.4679±.002 0.6127±.003 0.5196±.002 0.6200±.006
H2CGL 0.4579±.007 0.6812±.014 0.4337±.003 0.6573±.010 0.4727±.005 0.6836±.006
EGCN 0.8051±.009 0.1296±.009 0.6381±.007 0.2826±.020 0.7339±.012 0.2530±.021
Dysat 0.6297±.006 0.4242±.015 0.5412±.009 0.4705±.017 0.6207±.047 0.4354±.096
ROLAND 0.5978±.033 0.4911±.048 0.5437±.037 0.4495±.073 0.5521±.014 0.5718±.017
DisenGCN 0.8321±.010 0.1026±.030 0.6343±.006 0.2897±.016 0.6689±.011 0.3746±.018
DisenHAN 0.8221±.006 -0.0488±.038 0.8652±.181 -0.1110±.355 0.8766±.191 -0.1186±.552
CAL 0.7094±.018 0.1859±.222 0.5733±.015 0.4166±.026 0.5535±.009 0.5783±.013
DisC 0.6733±.006 0.3496±.013 0.5609±.031 0.4259±.065 0.5592±.021 0.5606±.035
DIDA 0.6203±.034 0.4449±.045 0.5108±.016 0.5354±.035 0.5723±.022 0.5328±.034
DyTed (EGCN) 0.8101±.008 0.1235±.008 0.6411±.007 0.2465±.006 0.7896±.009 0.1478±.025
DyTed (Dysat) 0.6820±.035 0.2426±.138 0.5386±.010 0.4602±.013 0.5838±.001 0.5029±.001
SILD 0.5836±.011 0.4977±.031 0.5009±.003 0.5629±.004 0.5448±.009 0.5889±.016
DPPDCC 0.4432±.005 0.6938±.008 0.4264±.002 0.6667±.005 0.4587±.002 0.6969±.002
#improve (%) 3.21 1.85 1.69 1.43 2.96 1.94
We run the experiments with three random seeds and report both the averaged results and standard deviations of all models. As shown in Figure 7, our proposed model DPPDCC consistently outperforms others with minimal variation, highlighting its capacity and stability.
Appendix F Complexity Experiment
Table 8. The input data, parameters, and inference time of selected baselines and DPPDCC.
Type model Input Data CS CHM PSY
Dynamic Hetero params inference params inference params inference
PLM SciBERT $\times$ $\times$ 111.1M 2.91s 111.1M 2.92s 111.1M 2.92s
Cascade CasSampling $\times$ $\times$ 0.2M 0.02s 0.2M 0.02s 0.2M 0.02s
Dynamic Dysat $\times$ 0.2M 0.66s 0.2M 0.76s 0.2M 1.24s
ROLAND 0.5M 4.76s 0.5M 6.59s 0.5M 8.71s
Disentangled CAL $\times$ $\times$ 0.3M 0.07s 0.3M 0.07s 0.3M 0.11s
DIDA $\times$ 1.4M 2.26s 1.4M 2.55s 1.4M 4.89s
DyTed $\times$ 2.1M 0.79s 2.1M 0.78s 2.1M 1.37s
SILD $\times$ 2.2M 1.31s 2.2M 1.42s 2.2M 2.40s
ours DPPDCC 14.3M 4.29s 31.0M 6.06s 14.3M 7.67s
Our model’s complexity is primarily influenced by the R-GCN architecture, which requires distinct message-passing. During training, proposed contrastive learning involves additional encoding of multi-view graphs, but this step can be omitted during inference. Notably, our model converges quickly, requiring only 5 epochs compared to 20 epochs for other baselines, ensuring comparable training time. Given that we train and test on datasets with approximately 0.2 million samples, it ensures comparable coverage. Figure 8 demonstrates that our model has comparable space occupancy and time complexity to the baseline models.
Appendix G Extended Visualizations
G.1. Weighting Score Visualization
(8.56/21.00)